find the five remaining trigonometric functions of theta, given that cos theta=-3/5 and sin theta > 0
Answer by lwsshak3(11628)

find the five remaining trigonometric functions of theta, given that cos theta=-3/5 and sin theta > 0
***
given data shows reference angle is in quadrant II
working with a (3-4-5)reference right triangle in quadrant II in which cos<0, sin>0
cos(theta)=-3/5
sin(theta)=4/5
tan(theta)=-4/3
cot(theta)=-3/4
sec(theta)=-5/3
csc(theta)=5/4